# Distribution-free data-driven smooth tests without $\chi^{2}$

### v2

@@ Abstract
This article demonstrates how recent developments in the theory of empirical processes allow us to construct a new family of asymptotically distribution-free smooth test statistics. Their distribution-free property is preserved even when the parameters are estimated, model selection is performed, and the sample size is only moderately large. A computationally efficient alternative to the classical parametric bootstrap is also discussed.
@@ 1 Introduction
Let $X$ be a continuous random variable with cumulative distribution function (CDF) $Q$ and probability density function (PDF) $q$ with support $\mathcal{X}\subseteq\mathbb{R}$. We are interested in assessing whether the unknown distribution $Q$ belongs to a family of distribution functions $G_{\bm{\beta}}$, with PDF $g_{\bm{\beta}}$, where $\bm{\beta}\in\mathcal{B}\subseteq\mathbb{R}^{p}$. To tackle this problem, we may consider an alternative model of the form
@@
where $\mathds{1}_{\{\cdot\}}$ denotes the indicator function. Clearly, the coefficients of the expansion in (1) are $\theta_{j\bm{\beta}}=\int_{\mathcal{X}}h_{j\bm{\beta}}(x)dQ(x)$, for all $j=1,\dots,m$.
@@
One could also rely on an exponential tilt to express the deviations of the alternative from $g_{\bm{\beta}}$ through the orthonormal set $\{h_{j\bm{\beta}}\}_{j=1}^{m}$, as in the original formulation of smooth tests proposed by Neyman, (1937). Other formulations may involve normalizing the expansion in (1) to ensure that the resulting density is non-negative (e.g., Gajek,, 1986). Here, we rely on the alternative in (1), proposed by Barton, (1953), by virtue of its simplicity. Nevertheless, the methods described in what follows can be adapted to other specifications of the alternative model for which the testing problem can be specified as in (2).
@@
Several test statistics have been proposed in the literature to test (2). Prominent examples include the score test statistic (Neyman,, 1937; Barton,, 1953), which consists of the sum of squares of the estimated coefficients in expansion (1) and typically employed for testing simple hypotheses; the generalized score test statistic (Thomas and Pierce,, 1979; Boos,, 1992) used for testing parametric hypotheses; and the order selection test statistic (Aerts et al.,, 1999) which naturally accounts for the selection of $m$ in (1). Unfortunately, while their asymptotic null distributions are known, in practical applications, a very large sample size is typically required to reach such limits. For instance, even when $G_{\bm{\beta}}$ is the normal distribution and $m$ in (1) is fixed, the generalized score statistic requires a sample size on the order of $\sim 10^{4}$ to be reasonably approximated by its asymptotic $\chi^{2}$ distribution (Klar,, 2000). Moreover, which basis functions to be included in the expansion in (1) is typically determined by data-driven selection criteria such as the Akaike information criterion (AIC) or Bayesian information criterion (BIC) (e.g., Ledwina,, 1994; Kallenberg and Ledwina,, 1995, 1997; Inglot et al.,, 1997; Inglot and Ledwina,, 2006). In this case, the selection process introduces additional sources of variability, which affect the limiting distributions of the test statistics.
To overcome these limitations, the parametric bootstrap is often employed by practitioners to derive the null distribution of the statistics used to test (2). However, such a numerical approach may become impractical when sampling from complicated distributions and/or when the estimation of the parameters is burdensome — as it is often the case in physics and astronomy (e.g., Balázs et al.,, 2017; Cusin et al.,, 2018, 2019). Moreover, the computational cost increases substantially when testing different models since the simulations must be performed on a case-by-case basis.
@@
This article demonstrates that most of the shortcomings arising when deriving the distributions of smooth test statistics, either asymptotically or numerically, can be overcome on the basis of recent developments in the theory of empirical processes. In particular, it is shown that the computational efficiency of the parametric bootstrap can be improved and its accuracy preserved by relying on the projected bootstrap (Cf. Section 3). Specifically, when the estimator of $\bm{\beta}$ is locally asymptotically linear, one can avoid recomputing it on each bootstrap replicate by relying on a projection of the underlying empirical process. The efficacy of this approach when deriving the distribution of classical goodness-of-fit statistics estimated via maximum likelihood has been investigated numerically by Algeri, (2022). Here, a formal proof is provided to confirm the validity of this approach in the context of smooth tests and for a more general class of estimators.
The projected bootstrap does not address scenarios in which numerical simulations are impractical due to the model’s complexity or the need to test multiple models simultaneously. Nevertheless, this limitation can be overcome by using the so-called Khmaladze-2 (K2) transform introduced by Khmaladze, (2016). In particular, we show that, in the context of smooth tests, such a transformation is especially valuable in that it enables the construction of a new set of basis functions, to be used in (1), which naturally lead to an entire new family of asymptotically distribution-free test statistics for (2). A suite of numerical studies demonstrates that the theoretical limiting null distribution of such tests well approximates the true null distribution even for samples that are only moderately large – i.e., of the order of $\sim 10^{2}$.
@@ 2 Smooth tests and the function-parametric empirical process
Let $X_{1},\ldots,X_{n}$ be independent and identically distributed (IID) random variables with CDF $Q$. Consider the Hilbert space $L^{2}(G_{\bm{\beta}})=\{h_{\bm{\beta}}:\left\langle h_{\bm{\beta}},h_{\bm{\beta}}\right\rangle_{G_{\bm{\beta}}}<\infty\},$ with inner product:
$\left\langle h_{\bm{\beta}},h^{\prime}_{\bm{\beta}}\right\rangle_{G_{\bm{\beta}}}=\int_{\mathcal{X}}h_{\bm{\beta}}(x)h^{\prime}_{\bm{\beta}}(x)dG_{\bm{\beta}}(x).$
For vector-valued pairs of functions $\bm{h}_{\bm{\beta}}=[h_{1\bm{\beta}},\dots,h_{t\bm{\beta}}]^{T},\bm{h}^{{}^{\prime}}_{\bm{\beta}}=[h^{\prime}_{1\bm{\beta}},\dots,h^{\prime}_{t^{\prime}\bm{\beta}}]^{T}$ in $L^{2}(G_{\bm{\beta}})$, denote $\big{\langle}{\bm{h}_{\bm{\beta}}^{T},\bm{h}^{{}^{\prime}T}_{\bm{\beta}}\big{\rangle}}_{G_{\bm{\beta}}}$ the outer product, which results in a $t\times t^{\prime}$ matrix, with $(i,j)$-thth element given by $\int h_{i\bm{\beta}}(x)h_{j\bm{\beta}}^{{}^{\prime}}(x)dG_{\bm{\beta}}(x).$ The function-parametric empirical process $v_{G,n}$ indexed by functions $h_{\bm{\beta}}$ in $L^{2}(G_{\bm{\beta}})$ is defined as:
$v_{G,n}(h_{\bm{\beta}})=\int_{\mathcal{X}}h_{\bm{\beta}}(x)dv_{\bm{\beta},n}(x)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}h_{\bm{\beta}}(X_{i})-\sqrt{n}\big{\langle}h_{\bm{\beta}},\bm{1}\big{\rangle}_{G_{\bm{\beta}}}$
where $\bm{1}$ denotes the function identically equal to one and
@@
is the classical empirical process. Hence, $v_{G,n}(h_{\bm{\beta}})=v_{\bm{\beta},n}(x)$ when $h_{\bm{\beta}}(z)=\mathds{1}_{\{z\leq x\}}$.
Most statistics proposed in the literature to test (2) can be specified as functionals of the process $v_{G,n}(h_{\bm{\beta}})$ and its projections. In particular, let
$\bm{u}_{\bm{\beta}}=\nabla_{\bm{\beta}}\ln g_{\bm{\beta}}\quad\text{and}\quad\Gamma_{\bm{\beta}}=\mathbb{E}_{G_{\bm{\beta}}}\left[-\nabla_{\bm{\beta}}\bm{u}^{T}_{\bm{\beta}}\right]$
be, respectively, the score function and the Fisher information matrix of $G_{\bm{\beta}}$. Define the orthonormalized score function in $L^{2}(G_{\bm{\beta}})$ as
$\bm{b}_{\bm{\beta}}=\Gamma_{\bm{\beta}}^{-1/2}\bm{u}_{\bm{\beta}}=\left[b_{\bm{\beta}_{1}},\ldots,b_{\bm{\beta}_{p}}\right]^{T},$
where $\Gamma_{\bm{\beta}}^{-1/2}$ denotes the principal square root matrix of $\Gamma_{\bm{\beta}}^{-1}$. Denote with $\{\widetilde{h}_{j\bm{\beta}}\}_{j=1}^{m}$ the “residuals” of $\{h_{j\bm{\beta}}\}_{j=1}^{m}$ after an orthogonal projection onto $\bm{b}_{\bm{\beta}}$, i.e.,
$\widetilde{h}_{j\bm{\beta}}=h_{j\bm{\beta}}-\bm{b}^{T}_{\bm{\beta}}\left\langle\bm{b}^{T}_{\bm{\beta}},h_{j\bm{\beta}}\right\rangle_{G_{\bm{\beta}}}=h_{j\bm{\beta}}-\sum_{k=1}^{p}b_{\bm{\beta}_{k}}\left\langle h_{j\bm{\beta}},b_{\bm{\beta}_{k}}\right\rangle_{G_{\bm{\beta}}},\text{ for }j=1,\ldots,m.$ (3)
[…]
$\widehat{S}_{m,n}=\sum_{i=1}^{m}\sum_{j=1}^{m}v_{G,n}(\widetilde{h}_{i\widehat{\bm{\beta}}_{n}})\bigl{(}\widehat{\Sigma}_{m}^{-1}\bigl{)}_{ij}v_{G,n}(\widetilde{h}_{j\widehat{\bm{\beta}}_{n}})$ (4)
where $\widehat{\Sigma}_{m}^{-1}$ is the inverse of estimated variance-covariance matrix, $\widehat{\Sigma}_{m}$, of elements
$\bigl{(}\widehat{\Sigma}_{m}\bigl{)}_{ij}=\mathbb{E}_{G_{\bm{\widehat{\beta}}_{n}}}\left[v_{G,n}(\widetilde{h}_{i\widehat{\bm{\beta}}_{n}})v_{G,n}(\widetilde{h}_{j\widehat{\bm{\beta}}_{n}})\right]$
and $\widehat{\bm{\beta}}_{n}$ denotes a locally asymptotically linear estimator of $\bm{\beta}$, defined as in Condition (A1) in Section 3.
When the number of basis functions, $m$, to be used in (1) and (4) is determined on the basis of the data observed, the so-called order selection test statistic (cfr. Aerts et al.,, 1999) can be used to incorporate the choice of the order $m$ directly into its formulation. Specifically, let $m$ be the maximizer of the selection criterion
@@
in which $\widehat{S}_{m,n}$ is the generalized score test statistic in (4) or its non-normalized counterpart, namely,
$\widehat{S}_{m,n}=\sum_{j=1}^{m}v^{2}_{G,n}(\widetilde{h}_{j\widehat{\bm{\beta}}_{n}}),$ (6)
with $\widehat{S}_{0,n}=0$ in both cases. $M_{n}$ denotes the maximum number of basis functions to be potentially considered, which could either be fixed or grow to infinity as $n\to\infty.$ Selecting $m$ using the criterion in (5) corresponds to choosing the first $m$ basis functions among $\{h_{j\bm{\beta}}\}_{j=1}^{\infty}$ to be included in the expansion (1).
The underlying idea of a test based on (5) is to reject the null hypothesis in (2) if the criterion in (5) is larger than zero for some $m$ in $\mathcal{M}_{n}=\left\{1,\ldots,M_{n}\right\}$; thus, an alternative model of the form in (1) is favored over $G_{\bm{\beta}}$. Equivalently, $G_{\bm{\beta}}$ is rejected when the order selection test statistic defined as
@@
is such that $\widehat{T}_{n}\geq C_{\alpha,n}$, for some constant $C_{\alpha,n}$ that controls the significance level, $\alpha$, of the test.
The criterion in (5) can be generalized so that basis functions indexed by any subset $B$ of $\mathcal{M}_{n}$, not necessarily ordered from $j=1,\dots,m$, are selected (Thas,, 2010, pp.103-107). In this case, the set of indices $B$ is chosen to maximize
@@
where $|B|$ denotes the cardinality of $B$ and $\widehat{S}_{B,n}$ is the counterpart of $\widehat{S}_{m,n}$ with the summations in (4) and (6) taken over the indexes in $B$. The corresponding subset selection test statistic $\widetilde{T}_{n}$ specifies as
@@
Compared to $\widehat{T}_{n}$ in (7), $\widetilde{T}_{n}$ offers greater flexibility in the selection of basis functions and thus accommodates a wider range of possible alternative models to be employed in the expansion in (1).
[…]

### v3

# A New Class of Asymptotically Distribution-Free Smooth Tests

@@ Abstract
This article demonstrates how recent developments in the theory of empirical processes allow us to construct a new family of asymptotically distribution-free smooth tests. Their distribution-free property is preserved even when the parameters are estimated, model selection is performed, and the sample size is only moderately large. A computationally efficient alternative to the classical parametric bootstrap is also discussed.
@@ 1 Introduction
Let $X$ be a continuous random variable with cumulative distribution function (CDF) $Q$ and probability density function (PDF) $q$ with support $\mathcal{X}\subseteq\mathbb{R}$. We are interested in assessing whether the unknown distribution $Q$ belongs to a family of distribution functions $G_{\bm{\beta}}$, with PDF $g_{\bm{\beta}}$, differentiable with respect to $\bm{\beta}\in\mathcal{B}\subseteq\mathbb{R}^{p}$. To tackle this problem, we may consider an alternative model of the form
@@
where $\mathds{1}_{\{\cdot\}}$ denotes the indicator function. In (1), as $m\rightarrow\infty$, the term in the curly brackets provides an orthonormal expansion for the density ratio $\frac{q}{g_{\bm{\beta}}}$. Hence, the coefficients of such an expansion correspond to
$\theta_{j\bm{\beta}}=\int_{\mathcal{X}}h_{j\bm{\beta}}(x)\frac{q(x)}{g_{\bm{\beta}}(x)}dG_{\bm{\beta}}(x)=\int_{\mathcal{X}}h_{j\bm{\beta}}(x)dQ(x),\quad\text{for all $j=1,\dots,m$}.$
@@
One could also rely on an exponential tilt to express the deviations of the alternative from $g_{\bm{\beta}}$ through the orthonormal set $\{h_{j\bm{\beta}}\}_{j=1}^{m}$, as in the original formulation of smooth tests proposed by Neyman (1937). Other formulations may involve normalizing the expansion in (1) to ensure that the resulting density is non-negative (e.g., Gajek, 1986). Here, we rely on the alternative in (1), proposed by Barton (1953), by virtue of its simplicity, and we implicitly assume that the $\theta_{j\bm{\beta}}$ coefficients are suitably constrained to ensure (1) is non-negative [^1] For example, as discussed in Barton (1953), when $h_{j\bm{\beta}}$ are the normalized shifted Legendre polynomials evaluated at the probability integral transform $G_{\bm{\beta}}(x)$ and $m=1$, non-negativity of (1) is achieved whenever $|\theta_{1\bm{\beta}}|<\frac{1}{3}$; whereas, if $m=2$, we must have $|\sqrt{5}\theta_{2\bm{\beta}}|\geq|\sqrt{3}\theta_{1\bm{\beta}}|$.. Nevertheless, the methods described in what follows can be adapted to other specifications of the alternative model for which the testing problem can be formulated as in (2).
@@
Several test statistics have been proposed in the literature for testing (2). Prominent examples include the score test statistic (Neyman, 1937; Barton, 1953), which consists of the sum of squares of the estimated coefficients in the expansion in (1), and typically employed for testing simple hypotheses; the generalized score test statistic [^2] Also known as the generalized smooth test statistic or efficient score test statistic. (Javitz, 1975; Thomas and Pierce, 1979) used for testing composite hypotheses; and the order selection test statistic (Aerts et al., 1999) which naturally accounts for the selection of $m$ in (1). Their asymptotic null distributions are known, but may require a very large sample size to reach such limits (Klar, 2000). Data-driven smooth tests (Ledwina, 1994) focus on the situation in which the basis functions to be included in the expansion in (1) are determined by data-driven selection criteria – such as BIC (e.g., Ledwina, 1994; Kallenberg and Ledwina, 1995, 1997; Inglot et al., 1997; Inglot and Ledwina, 2006b). Under the null model $G_{\bm{\beta}}$, classical smooth test statistics are asymptotically distribution-free.
For all the above-mentioned tests, asymptotic distribution-freeness arises as a consequence of the inherent structure of the test statistics, while the choice of basis $\{h_{j\bm{\beta}}\}_{j=1}^{\infty}$ remains somewhat arbitrary. We propose a shift of perspective in which a careful construction of the basis ensures that any test statistic based on its elements is asymptotically distribution-free by design. Such a construction relies on the Khmaladze-2 (K2) transform [^3] The name “Khmaladze-2 transform” is used to distinguish the transformation employed in this manuscript and proposed in Khmaladze (2016) from an earlier “Khmaladze transform” introduced by the same author in Khmaladze (1982). introduced by Khmaladze (2016), which enables both accurate and computationally efficient inference.
The K2 transform uses unitary operators to map model-dependent projections into a “standard projection” that is arbitrarily chosen by the user. This standard projection does not vary with the model being tested and is asymptotically distribution-free under the null hypothesis. In just over a decade, the K2 transformation has enabled several advances in the goodness-of-fit literature across various contexts—from testing discrete and multivariate parametric distributions (Khmaladze, 2013, 2016) to regression (Khmaladze, 2021, 2017) and grouped data analysis (Algeri and Khmaladze, 2026). By enabling its use in the context of smooth tests, this work extends its usefulness beyond goodness-of-fit. Specifically, it provides applied scientists with a new suite of distribution-free tests that offer a middle ground between goodness-of-fit tests and directional tests such as Neyman-Pearson.
@@
While asymptotically distribution-freeness is especially desirable when testing complex distributions and/or when computationally intensive parameter estimation is needed — as is often the case in physics and astronomy (e.g., Balázs et al., 2017; Cusin et al., 2018, 2019; Zhang et al., 2025; Algeri et al., 2025) — in other settings, applied scientists may choose to derive the null distributions of the test statistics by means of resampling methods. In Section 3, we discuss a computationally efficient alternative to the classical parametric bootstrap, called the projected bootstrap. Specifically, one can avoid re-estimating the unknown parameter $\bm{\beta}$ on each bootstrap replicate by exploiting a projection of the underlying empirical process. The efficacy of this approach when deriving the distribution of classical goodness-of-fit statistics estimated via maximum likelihood has been investigated numerically by Algeri (2022). Here, a formal proof is provided to confirm the validity of this approach in the context of smooth tests and for general $Z$-estimators.
@@ 2 Smooth tests and the function-parametric empirical process
Let $X_{1},\ldots,X_{n}$ be independent and identically distributed (IID) random variables with CDF $Q$. Consider the Hilbert space $L^{2}(G_{\bm{\beta}})=\{h:\left\langle h,h\right\rangle_{G_{\bm{\beta}}}<\infty\},$ with inner product:
$\left\langle h,h^{\prime}\right\rangle_{G_{\bm{\beta}}}=\int_{\mathcal{X}}h(x)h^{\prime}(x)dG_{\bm{\beta}}(x).$
The function-parametric empirical process $v_{\bm{\beta},n}(h)$ indexed by function $h$ in $L^{2}(G_{\bm{\beta}})$ is defined as:
$v_{\bm{\beta},n}(h)=\int_{\mathcal{X}}h(x)dv_{\bm{\beta},n}(x)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\bigl[h(X_{i})-\mathbb{E}_{G_{\bm{{\beta}}}}[h(X_{i})]\bigl]$
and
@@
is the classical empirical process. Hence, $v_{\bm{\beta},n}(h)=v_{\bm{\beta},n}(x)$ when $h(z)=\mathds{1}_{\{z\leq x\}}$.
Most statistics proposed in the literature to test (2) can be specified as functionals of the process $v_{\bm{\beta},n}(h)$ and its projections. In particular, let
$\bm{u}_{\bm{\beta}}=\nabla_{\bm{\beta}}\ln g_{\bm{\beta}}\quad\text{and}\quad\Gamma_{\bm{\beta}}=\left\langle\bm{u}_{\bm{\beta}},\bm{u}^{T}_{\bm{\beta}}\right\rangle_{G_{\beta}}$
be, respectively, the score function and the Fisher information matrix of $G_{\bm{\beta}}$. Define the orthonormalized score function, with coordinates in $L^{2}(G_{\bm{\beta}})$, as
$\bm{b}_{\bm{\beta}}=\left[b_{\bm{\beta}_{1}},\ldots,b_{\bm{\beta}_{p}}\right]^{T}=\Gamma_{\bm{\beta}}^{-1/2}\bm{u}_{\bm{\beta}},$
where $\Gamma_{\bm{\beta}}^{-1/2}$ denotes the principal square root matrix of $\Gamma_{\bm{\beta}}^{-1}$. Denote with $\{\widetilde{h}_{j\bm{\beta}}\}_{j=1}^{m}$ the “residuals” of $\{h_{j\bm{\beta}}\}_{j=1}^{m}$ in (1) after an orthogonal projection onto $\bm{b}_{\bm{\beta}}$, i.e.,
$\widetilde{h}_{j\bm{\beta}}=h_{j\bm{\beta}}-\bm{b}^{T}_{\bm{\beta}}\left\langle\bm{b}_{\bm{\beta}},h_{j\bm{\beta}}\right\rangle_{G_{\bm{\beta}}}=h_{j\bm{\beta}}-\sum_{k=1}^{p}b_{\bm{\beta}_{k}}\left\langle h_{j\bm{\beta}},b_{\bm{\beta}_{k}}\right\rangle_{G_{\bm{\beta}}},\text{ for }j=1,\ldots,m.$ (3)
Deviations within the parametric family $\{G_{\bm{\beta}},\ \bm{\beta}\in\mathcal{B}\}$ occur in the direction given by the the score function $\bm{u}_{\bm{\beta}}$. Therefore, since the functions $\widetilde{h}_{j\bm{\beta}}$ are obtained by projecting $h_{j\bm{\beta}}$ onto the orthogonal complement of the tangent space given by the span of $\bm{u}_{\bm{\beta}}$, they capture deviations outside the hypothesized parametric family.
The most widely used statistics for testing (2) is the generalized score test statistic, (cfr. Javitz, 1975; Thomas and Pierce, 1979), defined as
$\widehat{S}_{m,n}=\sum_{i=1}^{m}\sum_{j=1}^{m}v_{\widehat{\bm{\beta}}_{n},n}(\widetilde{h}_{i\widehat{\bm{\beta}}_{n}})\bigl(\Sigma^{-1}_{m,\widehat{\bm{\beta}}_{n}}\bigl)_{ij}v_{\widehat{\bm{\beta}}_{n},n}(\widetilde{h}_{j\widehat{\bm{\beta}}_{n}})$ (4)
where $\Sigma^{-1}_{m,\widehat{\bm{\beta}}_{n}}$ is the inverse of estimated variance-covariance matrix, ${\Sigma}_{m,\widehat{\bm{\beta}}_{n}}$, of elements
$\bigl({\Sigma}_{m,\widehat{\bm{\beta}}_{n}}\bigl)_{ij}=\mathbb{E}_{G_{\bm{{\beta}}}}\left[v_{\bm{\beta},n}(\widetilde{h}_{i{\bm{\beta}}})v_{\bm{\beta},n}(\widetilde{h}_{j{\bm{\beta}}})\right]\Bigl|_{\bm{\beta}=\bm{\widehat{\beta}}_{n}},$
which we assume to be a continuous function in $\bm{\beta}$, and $\widehat{\bm{\beta}}_{n}$ denotes a consistent estimator of $\bm{\beta}$ (see condition (A1) in Section 3).
When $m$, the number of basis functions to be used in (1) and (4), is determined on the basis of the data observed, the so-called order selection test statistic (cfr. Aerts et al., 1999) can be used to incorporate the choice of the order $m$ directly into its formulation. Specifically, let $m$ be the maximizer of the selection criterion
@@
in which $\widehat{S}_{m,n}$ is the generalized score test statistic in (4) or its non-normalized counterpart:
$\widehat{S}_{m,n}=\sum_{j=1}^{m}v^{2}_{\widehat{\bm{\beta}}_{n},n}(\widetilde{h}_{j\widehat{\bm{\beta}}_{n}}).$ (6)
Let $M_{n}$ denote the maximum number of basis functions to be considered, which could either be fixed or grow to infinity as $n\to\infty.$ Selecting $m$ using the criterion in (5) corresponds to choosing the first $m$ basis functions among $\{h_{j\bm{\beta}}\}_{j=1}^{M_{n}}$ to be included in the expansion in (1).
The underlying idea of a test based on (5) is to reject the null hypothesis in (2) if the criterion in (5) is larger than zero for some $m$ in $\mathcal{M}_{n}=\left\{1,\ldots,M_{n}\right\}$ – that is, when an alternative model of the form in (1) is favored over $G_{\bm{\beta}}$. Equivalently, $G_{\bm{\beta}}$ is rejected when the order selection test statistic defined as
@@
is such that $\widehat{T}_{n}\geq C_{\alpha,n}$, for some constant $C_{\alpha,n}$ that controls the significance level $\alpha$ of the test.
The criterion in (5) can be generalized so that basis functions with indexes $j$ in any subset $B$ of $\{1,\dots,M_{n}\}$, not necessarily ordered from $j=1,\dots,m$, are selected (Thas, 2010, pp.103-107). In this case, the set of indices $B$ is chosen to maximize
@@
where $|B|$ denotes the cardinality of $B$ and $\widehat{S}_{B,n}$ is the counterpart of $\widehat{S}_{m,n}$ with the summations in (4) and (6) taken over the indexes in $B$. The corresponding subset selection test statistic specifies as
@@
Compared to $\widehat{T}_{n}$ in (7), $\widetilde{T}_{n}$ offers greater flexibility in the selection of the basis functions and thus accommodates a wider range of possible alternative models to be employed in the expansion in (1).
The limiting distribution of the subset selection statistic in (9) cannot be easily derived. Therefore, a computationally efficient algorithm is required to simulate its null distribution, especially when $M_{n}$ is large. In the next section, we introduce an alternative to the classical parametric bootstrap to perform this task.